# Elimination Distance to Dominated Clusters

### v1

@@ Abstract
In the Dominated Cluster Deletion problem we are given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether there exists a set of at most $k$ vertices whose removal results in a graph in which each connected component has a dominating set of size at most $d$. In the Elimination Distance to Dominated Clusters problem we are again given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether we can recursively delete vertices up to depth $k$ such that each remaining connected component has a dominating set of size at most $d$. Bentert et al. [Bentert et al., MFCS 2024] recently provided an almost complete classification of the parameterized complexity of Dominated Cluster Deletion with respect to the parameters $k$, $d$, $c$, and $\Delta$, where $c$ and $\Delta$ are the degeneracy, and the maximum degree of the input graph, respectively. In particular, they provided a non-uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$. They left as an open problem whether the problem is fixed-parameter tractable with respect to the parameter $k+d+c$. We provide a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters. We furthermore show that both problems are FPT when parameterized by $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph, a parameter that is upper bounded and may be much smaller than the degeneracy $c$, positively answering the open question of Bentert et al. We almost complete the picture by providing an almost full classification for the parameterized complexity and kernelization complexity of Elimination Distance to Dominated Clusters. The one difficult base case that remains open is whether Treedepth (the case $d=0$) is NP-hard on graphs of bounded maximum degree.
@@ 1 Introduction
[…]
@@
A related measure is the elimination distance to a class $\mathscr{C}$, which measures the number of recursive deletions of vertices needed for a graph $G$ to become a member of $\mathscr{C}$. Formally, a graph $G$ has elimination distance 00 to $\mathscr{C}$ if $G\in\mathscr{C}$, and otherwise elimination distance $k+1$, if in every connected component of $G$, we can delete a vertex such that the resulting graph has elimination distance $k$ to $\mathscr{C}$. Note that elimination distance to the class containing only the empty graph corresponds to the well-known measure treedepth. Elimination distance was introduced by Bulian and Dawar [9] in their study of the parameterized complexity of the graph isomorphism problem. Just as distance to triviality, but with more general applicability, elimination distance allows to lift algorithmic results from a base class to more general graph classes. For example, Bulian and Dawar [9] provided an FPT algorithm for the graph isomorphism problem parameterized by the elimination distance to the class $\mathscr{C}_{d}$ of graphs with maximum degree bounded by $d$, for any fixed integer $d$. Hols et al. [27] proved the existence of polynomial kernels for the vertex cover problem parameterized by the size of a deletion set to graphs of bounded elimination distance to different classes of graphs. Agrawal et al. [1] proved fixed-parameter tractability of Elimination Distance to $\mathscr{C}$ for any class $\mathscr{C}$ that can be defined by a finite set of forbidden induced subgraphs. Agrawal and Ramanujan studied the Elimination Distance to Cluster Graphs [2]. Fomin, Golovach and Thilikos further generalized several of these prior results by considering elimination distances to graph classes expressible by restricted first-order formulas [22]. These results are also implied by the recent algorithmic meta theorem for separator logic [31, 33]. The related concept of recursive backdoors has recently been studied in the context of efficient SAT solving [17, 30].
@@
Our main contribution are the uniform FPT algorithms for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on semi-ladder-free graphs.
@@
[…]
@@ 1.1 Techniques
Lokshtanov et al. [29] proved a powerful meta theorem for logically defined graph properties. Whenever a CMSO-property (monadic second-order logic with modulo counting predicates) can be solved efficiently on unbreakable graphs (for appropriately chosen parameters $k$ and $q$), then it can also be solved efficiently on general graphs. There is a small caveat though. The meta theorem uses the non-constructive recursive understanding technique, and hence, one only obtains an efficient non-uniform algorithm on general graphs. Both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are CMSO-properties, and hence fall into this framework. Bentert et al. rely on this framework and show how to solve Dominated Cluster Deletion on unbreakable graphs in time $f(k,d)\cdot n^{\mathcal{O}(d)}$. The key combinatorial observation is that in $(q,k)$-unbreakable graphs, by the deletion of at most $k$ vertices leads to one giant component and a bounded number of remaining small components. One can then guess a dominating set for the giant component in time $\mathcal{O}(n^{d})$ and argue combinatorially to efficiently find the vertices to be deleted. Using the result of Lokshtanov et al. one obtains the non-uniform algorithm for general graphs with the desired running time.
We remark that Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are in fact a first-order definable problems, that is, for all parameters $k,d$ there exist first-order formulas of length $\mathcal{O}(k,d)$ defining the problems. Hence, the problems are fixed-parameter tractable on all classes on which first-order model checking is fixed-parameter tractable, e.g. on nowhere dense classes [24] and even monadically stable graph classes [15, 16], as well as on classes with bounded cliquewidth [10] (and classes with bounded twinwidth [8] when contraction sequences are given with the input). Note that first-order logic cannot define connected components in general, however, the connected components arising in the recursive deletion of elements must have small diameter when they can be dominated by few vertices, hence, connected components in positive instances of Dominated Cluster Deletion and Elimination Distance to Dominated Clusters become first-order definable. Thus, on these classes one obtains algorithms with running time $f(k,d)\cdot n^{\mathcal{O}(1)}$ by the meta theorems for first-order logic. It is conjectured that monadically dependent classes are the most general hereditary graph classes that admit efficient first-order model checking, we hence included them in the inclusion diagram below.
@@
As mentioned, we follow the approach of Bentert et al. [3] to reduce to unbreakable graphs. We first observe that on unbreakable graphs, the notion of elimination distance almost collapses to distance to triviality, which allows us to treat Dominated Cluster Deletion and Elimination Distance to Dominated Clusters in a similar way. In order to avoid the non-uniform approach of Lokshtanov et al., we use a decomposition theorem due to Cygan et al. [12]. The theorem states that every graph admits a tree decomposition with small adhesion such that each bag is unbreakable in the subgraph induced by the bags below. We show that if an annotated version of the Partial Dominating Set problem can be solved in a slightly modified graph (that we call the bag graph), then we can implement an efficient dynamic programming algorithm along the tree decomposition to the original input graph for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters.
The (Annotated) Partial Dominating Set problem is the following problem. We are given a graph with three sets $F$, $R$ and $B$ and parameters $d$ and $k$. We ask whether there exist $k$ vertices of $G-F$ ($F$ stands for forbidden) that may be deleted such that the remaining graph has a Red-Blue dominating set of size at most $d$, i.e., the vertices of $R$ can be dominated by $d$ vertices of $B$ ($R$ and $B$ stand for Red and Blue). Formally, the problem is defined as follows. Given a graph $G$ with unary predicates $F\subseteq V(G)$, $R\subseteq V(G)$, and $B\subseteq V(G)$ and parameters $d$ and $k$, does there exist a set $D\subseteq B$ of size at most $d$ and a set $X\subseteq V(G)-F$ of size at most $k$ such that $G[R-X]$ is dominated by $D$? The problem can obviously be solved in time $n^{\mathcal{O}(d)}$ (guess the dominating set and check if at most $k$ Red vertices remain undominated). By our dynamic programming approach we hence get a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ in general graphs. We then turn our attention to restricted graph classes where (Annotated) Partial Domination can be solved more efficiently. Using an approach of Fabiański et al. [19] we observe that (Annotated) Partial Domination can be solved efficiently on all semi-ladder-free graph classes, which yields our second main result.
@@
As mentioned, we require that (Annotated) Partial Domination needs to be solved on a slightly modified graph class (on the bag graphs), however, the property of being semi-ladder-free is preserved by this modification, so that ultimately we conclude our main results, Theorem 1.1 and Theorem 1.2.
Organization. The paper is organized as follows. After recalling the necessary notation in Section 2, we talk more precisely about partial domination and the graph parameters enabling FPT evaluation in Section 3. We then present how to solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graph with bounded semi-ladder index in Section 4. After that, in Section 5, we sketch the uniform algorithm on general semi-ladder-free graphs for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$. The details are in appendix. The hardness results conclude the paper in Section 6.
@@ 2 Preliminaries
Graphs. We consider finite, undirected graphs without loops. Our notation is standard, and we refer to Diestel’s textbook for more background on graphs [13]. We write $|G|$ for the size of the vertex set $V(G)$ and $\|G\|$ for $|V(G)|+|E(G)|$, where $E(G)$ is the edge set of a graph $G$. A connected component of $G$ is a maximal connected subgraph of $G$. For a vertex subset $X\subseteq V(G)$ we write $G[X]$ for the subgraph of $G$ induced by $X$. We write $G-X$ for $G[V(G)\setminus X]$ and for singleton sets $\{v\}$ we write $G-v$ instead of $G-\{v\}$. We denote the open neighborhood of a vertex $v$ by $N(v)$ and the closed neighborhood (including the vertex $v$) by $N[v]$. For a set $Y\subseteq V(G)$ we write $N[Y]$ for $\bigcup_{v\in Y}N[y]$.
Semi-ladders. Let $G$ be a graph. Two sequences, $a_{1},\ldots,a_{n}\in V(G)$ and $b_{1},\ldots,b_{n}\in V(G)$ of $2n$ distinct vertices, form a semi-ladder of order $n$ in $G$ if $\{a_{i},b_{j}\}\in E(G)$ for all $i,j\in\{1,\ldots,n\}$ with $i>j$, and $\{a_{i},b_{i}\}\notin E(G)$ for all $i\in\{1,\ldots,n\}$. Note that we do not impose any condition for $i<j$. The semi-ladder index of a graph is the maximum order of a semi-ladder that it contains. A class of graphs has bounded semi-ladder index or is called semi-ladder-free if there is some $\ell$ such that the semi-ladder index of all of its members is bounded by $\ell$, see Figure 2.
@@
  • For every node $x\in V(T)$, the bag $\mathsf{bag}(x)$ is $(q,k)$-unbreakable in the subgraph induced by $G[\mathsf{cone}(x)]$.
@@
We may further assume that the tree decomposition we get from Theorem 2.1 is regular, see the construction of Lemma 2.8 of [7] and the discussion in [31].
@@ 3 Partial domination and FPT evaluation
We show how to efficiently solve Annotated Partial Domination problem on semi-ladder-free graphs. This problem falls into the framework of domination type problems on semi-ladder-free graphs introduced by Fabiański et al. [19].
We are going to use the following meta theorem of Fabiański et al. [19]. A domination-type problem is a formula $\exists\bar{x}\forall\bar{y}\delta(\bar{x},\bar{y})$, where $\delta$ is a positive first-order formula using fixed distances as predicates (called a distance formula in [19]). On every graph $G$ a formula $\delta(\bar{x},\bar{y})$ defines a bipartite graph $G_{\delta}$ on the vertex $V(G)^{|\bar{x}|}\times V(G)^{|\bar{y}|}$, where two elements $\bar{u}$ and $\bar{v}$ are connected by an edge if the formula $\delta(\bar{u},\bar{v})$ is true in $G$. The semi-ladder index of $\delta$ on $G$ is defined as the semi-ladder index of $G_{\delta}$.
@@
Let $\mathscr{C}$ be a class of graphs and let $\delta$ be a distance formula. Assume that $G_{\delta}$ has bounded semi-ladder index $\ell$ for some fixed $\ell\in\mathbb{N}$ on all graphs $G\in\mathscr{C}$. Then there is an algorithm that solves the domination-type problem $\delta$ on graphs $G$ from $\mathscr{C}$ in time $f(\ell,|\varphi|)\cdot\|G\|$.
Annotated Partial Domination can be written with the following first order formula:
$\varphi(k,d)=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}\neg F(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
@@
However, $\varphi$ is not a distance formula as it uses negations. This can be circumvented by inverting the predicate $F^{\prime}=V(G)-F$ and $R^{\prime}=V(G)-R$. We then obtain the following formula
@@
While the semi-ladder index is currently one of the most general graph parameter enabling FPT evaluation of $\delta_{k,d}$, some other parameter could have this property.
@@
A graph parameter $t(\cdot)$ has 1 if for every graph $G$ Annotated Partial Domination can be solved in time $f(d,k,t(G))\cdot|G|^{\mathcal{O}(1)}$ for a computable function $f$.
As previously discussed, examples of parameters with 1 include treewidth, maximum degree, Hadwiger number (i.e. largest clique minor), degeneracy, largest excluded bi-clique, as they all imply bounded semi-ladder index. Other parameters, which are however not as interesting, as the problem can be solved by efficient FO model checking, are cliquewidth and twinwidth (when contraction sequences are given with the input).
As mentioned, we want to solve Annotated Partial Domination on slightly modified graphs. See Definition B.6 for the precise definition of bag graphs. We therefore want a parameter that, if bounded on the graph, bounds a parameter with 1 on bag graphs.
@@
A graph parameter $t_{2}(\cdot)$ has 2 if there is a parameter $t_{1}(\cdot)$ satisfying 1 such that for every graph $G$, and integers $q,k,d$, every $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$, vertex $x\in T$, and set $S\subseteq\mathsf{cone}(x)$ we have that $t_{1}(\mathsf{bgraph}_{S}(x))\leq f(q,k,d,t_{2}(G))$.
In the rest of the paper, we precisely reference parameters having 1 or 2. In particular, the semi-ladder index also satisfies 2, see Lemma B.11.

### v2

@@ Abstract
In the Dominated Cluster Deletion problem, we are given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether there exists a set of at most $k$ vertices whose removal results in a graph in which each connected component has a dominating set of size at most $d$. In the Elimination Distance to Dominated Clusters problem, we are again given an undirected graph $G$ and integers $k$ and $d$ and the question is to decide whether we can recursively delete vertices up to depth $k$ such that each remaining connected component has a dominating set of size at most $d$. Bentert et al. [Bentert et al., MFCS 2024] recently provided an almost complete classification of the parameterized complexity of Dominated Cluster Deletion with respect to the parameters $k$, $d$, $c$, and $\Delta$, where $c$ and $\Delta$ are the degeneracy, and the maximum degree of the input graph, respectively. In particular, they provided a non-uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$. They left as an open problem whether the problem is fixed-parameter tractable with respect to the parameter $k+d+c$. We provide a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters. We furthermore show that both problems are FPT when parameterized by $k+d+\ell$, where $\ell$ is the semi-ladder index of the input graph, a parameter that is upper bounded and may be much smaller than the degeneracy $c$, positively answering the open question of Bentert et al. We further complete the picture by providing an almost full classification for the parameterized complexity and kernelization complexity of Elimination Distance to Dominated Clusters. The one difficult base case that remains open is whether Treedepth (the case $d=0$) is NP-hard on graphs of bounded maximum degree.
@@ 1 Introduction
Assume that $\mathcal{Q}$ is a graph problem that is hard to solve on general graphs but $\mathscr{C}$ is a class of graphs where $\mathcal{Q}$ can be solved efficiently. In many cases, we can also solve $\mathcal{Q}$ efficiently on instances that are close to the instances of $\mathscr{C}$, say, on graphs that belong to $\mathscr{C}$ after the deletion of a small number of vertices. Guo et al. [25] formalized this concept under the name distance from triviality. For example, the size of a minimum vertex cover is the distance to the class of edgeless graphs and the size of a minimum feedback vertex set is the distance to the class of acyclic graphs. Cluster Vertex Deletion is the problem to determine for a given graph and number $k$ whether there exists a set of at most $k$ vertices whose deletion results in a cluster graph, that is, a graph in which every connected component is a clique [5, 34]. Many generalizations and special cases of Cluster Vertex Deletion have been studied. For example, one may require that each connected component in the resulting graph is an $s$-club (a graph of diameter at most s) [21, 28] or an $s$-plex (a graph in which each vertex has degree at least nsn-s) [26, 35]. Note that the special case where each clique in the solution graph has to be of size one is again the Vertex Cover problem. Very recently, Bentert et al. [4] studied the following Dominated Cluster Deletion problem: Given a graph $G$ and integer parameters $k$ and $d$, does there exist a set $S$ of at most $k$ vertices such that each connected component of GSG-S can be dominated by at most $d$ vertices? Bentert et al. provided an almost complete classification of the parameterized complexity and kernelization complexity of Dominated Cluster Deletion with respect to the parameters $k$, $d$, $c$, and $\Delta$, where $c$ and $\Delta$ are the degeneracy and the maximum degree of the input graph, respectively. They proved the following results: Dominated Cluster Deletion is para-NP-hard for parameters $k+\Delta$ and $d+\Delta$, $W[2]$-hard for parameter $k+d$ and admits a non-uniform algorithm with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$, and is non-uniformly fixed-parameter tractable with respect to parameter $k+d+\Delta$. They left as an open problem whether the problem is fixed-parameter tractable with respect to the parameter $k+d+c$. They also showed that the problem does not admit a polynomial kernel even for $d=1$ with respect to the parameter $k+c$, or with respect to the parameter $k+d+\Delta$, unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
@@
A related measure is the elimination distance to a class $\mathscr{C}$, which measures the number of recursive deletions of vertices needed for a graph $G$ to become a member of $\mathscr{C}$. Formally, a graph $G$ has elimination distance 0 to $\mathscr{C}$ if $G\in\mathscr{C}$, and otherwise elimination distance $k+1$, if in every connected component of $G$, we can delete a vertex such that the resulting graph has elimination distance $k$ to $\mathscr{C}$. Note that elimination distance to the class containing only the empty graph corresponds to the well-known measure treedepth. Elimination distance was introduced by Bulian and Dawar [10] in their study of the parameterized complexity of the graph isomorphism problem. Just as distance to triviality, but with more general applicability, elimination distance allows to lift algorithmic results from a base class to more general graph classes. For example, Bulian and Dawar [10] provided an FPT algorithm for the graph isomorphism problem parameterized by the elimination distance to the class $\mathscr{C}_{d}$ of graphs with maximum degree bounded by $d$, for any fixed integer $d$. Hols et al. [27] proved the existence of polynomial kernels for the vertex cover problem parameterized by the size of a deletion set to graphs of bounded elimination distance to different classes of graphs. Agrawal et al. [2] proved fixed-parameter tractability of Elimination Distance to $\mathscr{C}$ for any class $\mathscr{C}$ that can be defined by a finite set of forbidden induced subgraphs. Agrawal and Ramanujan studied the Elimination Distance to Cluster Graphs [3]. Fomin, Golovach and Thilikos further generalized several of these prior results by considering elimination distances to graph classes expressible by restricted first-order formulas [22]. These results are also implied by the recent algorithmic meta theorem for separator logic [31, 33]. The related concept of recursive backdoors has recently been studied in the context of efficient SAT solving [18, 30]. To the best of our knowledge, we are the first to study elimination distance to classes with small dominating sets.
@@
Our main contributions are the uniform FPT algorithms for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on semi-ladder-free graphs.
@@
    The problem does not admit a polynomial kernel with respect to parameter $k$ even for fixed parameters $d=0$ and $c=2$ unless $\textsf{NP}\subseteq\textsf{coNP/poly}$; The parameter $d=0$ corresponds to Treedepth. We prove the simple observation that Treedepth is NP-complete on $2$-degenerate graphs. It is then trivial using an AND-cross-composition to prove that Treedepth on $2$-degenerate graphs does not admit a polynomial kernel unless $\textsf{NP}\subseteq\textsf{coNP/poly}$.
@@ 1.1 Techniques
Lokshtanov et al. [29] proved a powerful meta theorem for logically defined graph properties. Whenever a CMSO-property (monadic second-order logic with modulo counting predicates) can be solved efficiently on unbreakable graphs (for appropriately chosen parameters $k$ and $q$), then it can also be decided efficiently on general graphs. There is a small caveat though. The meta theorem uses the non-constructive recursive understanding technique, and hence, one only obtains an efficient non-uniform algorithm on general graphs. Both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are CMSO-properties, and hence fall into this framework. Bentert et al. rely on this framework and show how to solve Dominated Cluster Deletion on unbreakable graphs in time $f(k,d)\cdot n^{\mathcal{O}(d)}$. The key combinatorial observation is that in $(q,k)$-unbreakable graphs, the deletion of at most $k$ vertices leads to one giant component and a bounded number of remaining small components. One can then guess a dominating set for the giant component in time $\mathcal{O}(n^{d})$ and argue combinatorially to efficiently find the vertices to be deleted. Using the result of Lokshtanov et al. one obtains the non-uniform algorithm for general graphs with the desired running time.
We remark that Dominated Cluster Deletion and Elimination Distance to Dominated Clusters are in fact first-order definable problems, that is, for all parameters $k,d$, there exist first-order formulas of length $\mathcal{O}_{k,d}(1)$ defining the problems. Hence, the problems are fixed-parameter tractable on all classes on which first-order model checking is fixed-parameter tractable, e.g. on nowhere dense classes [24] and even monadically stable graph classes [16, 17], as well as on classes with bounded cliquewidth [11] (and classes with bounded twinwidth [9] when contraction sequences are given with the input). Note that first-order logic cannot define connected components in general. However, the connected components arising in the recursive deletion of elements must have small diameter when they can be dominated by few vertices. Hence, connected components in positive instances of Dominated Cluster Deletion and Elimination Distance to Dominated Clusters become first-order definable. Thus, on these classes one obtains algorithms with running time $f(k,d)\cdot n^{\mathcal{O}(1)}$ by the meta theorems for first-order logic. It is conjectured that monadically dependent classes are the most general hereditary graph classes that admit efficient first-order model checking, we hence included them in the inclusion diagram, see Figure˜1.
@@
As mentioned, we follow the approach of Bentert et al. [4] to reduce to unbreakable graphs. We first observe that on unbreakable graphs, the notion of elimination distance almost collapses to distance to triviality, enabling us to treat Dominated Cluster Deletion and Elimination Distance to Dominated Clusters in a similar way. This idea has already been examined for hereditary classes in several papers see e.g. [1]. Note that being a dominated cluster is not a hereditary notion. In order to avoid the non-uniform approach of Lokshtanov et al. [29], we use a decomposition theorem due to Cygan et al. [13]. The theorem states that every graph admits a tree decomposition with small adhesion such that each bag is unbreakable in the subgraph induced by the bags below. We show that if an annotated version of the Partial Dominating Set problem can be solved in a slightly modified graph (that we call the bag graph), then we can implement an efficient dynamic programming algorithm along the tree decomposition to the original input graph for both Dominated Cluster Deletion and Elimination Distance to Dominated Clusters.
The (Annotated) Partial Dominating Set problem is the following problem. We are given a graph with three sets $F$, $R$ and $B$ and parameters $d$ and $k$. We ask whether there exist $k$ vertices of GFG-F ($F$ stands for forbidden) that may be deleted such that the remaining graph has a Red-Blue dominating set of size at most $d$, i.e. the vertices of $R$ can be dominated by $d$ vertices of $B$ ($R$ and $B$ stand for Red and Blue). Formally, the problem is defined as follows. Given a graph $G$ with (not necessarily disjoint) sets $F\subseteq V(G)$, $R\subseteq V(G)$, and $B\subseteq V(G)$ and parameters $d$ and $k$, does there exist a set $X\subseteq V(G)-F$ of size at most $k$ and a set $D\subseteq B-X$ of size at most $d$ such that G[RX]G[R-X] is dominated by $D$? The problem can obviously be solved in time $n^{\mathcal{O}(d)}$ (guess the dominating set and check if at most $k$ Red vertices remain undominated). By our dynamic programming approach, we hence get a uniform algorithm running in time $f(k,d)\cdot n^{\mathcal{O}(d)}$ in general graphs. We then turn our attention to restricted graph classes where (Annotated) Partial Domination can be solved more efficiently. Using an approach of Fabiański et al. [20], we observe that (Annotated) Partial Domination can be solved efficiently on all semi-ladder-free graph classes, which yields our second main result.
@@
As mentioned, we require that (Annotated) Partial Domination needs to be solved on a slightly modified graph class (on the bag graphs), however, the property of being semi-ladder-free is preserved by this modification, so that ultimately we conclude our main results, Theorem˜1.1 and Theorem˜1.2.
Organization. The paper is organized as follows. After recalling the necessary notation in Section˜2, we talk more precisely about partial domination and the graph parameters enabling FPT evaluation in Section˜3. We then present how to solve Dominated Cluster Deletion and Elimination Distance to Dominated Clusters on unbreakable graph with bounded semi-ladder index in Section˜4. After that, in Section˜5, we sketch the uniform algorithm on general semi-ladder-free graphs for Dominated Cluster Deletion and Elimination Distance to Dominated Clusters with running time $f(k,d)\cdot n^{\mathcal{O}(d)}$. The details are in the appendix. The hardness results conclude the paper in Section˜6.
@@ 2 Preliminaries
Graphs. We consider finite, undirected graphs without loops. Our notation is standard, and we refer to Diestel’s textbook for more background on graphs [14]. We write $|G|$ for the size of the vertex set $V(G)$ and $\|G\|$ for $|V(G)|+|E(G)|$, where $E(G)$ is the edge set of a graph $G$. A connected component of $G$ is a maximal connected subgraph of $G$. For a vertex subset $X\subseteq V(G)$, we write $G[X]$ for the subgraph of $G$ induced by $X$. We write GXG-X for $G[V(G)\setminus X]$ and for singleton sets $\{v\}$ we write GvG-v instead of $G-\{v\}$. We denote the open neighborhood of a vertex $v$ by $N(v)$ and the closed neighborhood (including the vertex $v$) by $N[v]$. For a set $Y\subseteq V(G)$, we write $N[Y]$ for $\bigcup_{v\in Y}N[y]$.
Semi-ladders. Let $G$ be a graph. Two sequences, $a_{1},\ldots,a_{n}\in V(G)$ and $b_{1},\ldots,b_{n}\in V(G)$ of $2n$ distinct vertices, form a semi-ladder of order $n$ in $G$ if $\{a_{i},b_{j}\}\in E(G)$ for all $i,j\in\{1,\ldots,n\}$ with $i>j$, and $\{a_{i},b_{i}\}\notin E(G)$ for all $i\in\{1,\ldots,n\}$. Note that we do not impose any condition for $i<j$. The semi-ladder index of a graph is the maximum order of a semi-ladder that it contains. A class of graphs has bounded semi-ladder index or is called semi-ladder-free if there is some $\ell$ such that the semi-ladder index of all of its members is bounded by $\ell$, see Figure˜2.
@@
  • For every node $x\in V(T)$, the bag $\mathsf{bag}(x)$ is $(q,k)$-unbreakable in the subgraph $G[\mathsf{cone}(x)]$.
@@
We may further assume that the tree decomposition we get from Theorem˜2.1 is regular, see the construction of Lemma 2.8 of [8] and the discussion in [31].
@@ 3 Partial domination and FPT evaluation
We show how to efficiently solve the Annotated Partial Domination problem on semi-ladder-free graphs. This problem falls into the framework of domination-type problems on semi-ladder-free graphs introduced by Fabiański et al. [20].
We are going to use the following meta theorem of Fabiański et al. [20]. A domination-type problem is a formula $\exists\bar{x}\forall\bar{y}\delta(\bar{x},\bar{y})$, where $\delta$ is a positive first-order formula using fixed distances as predicates (called a distance formula in [20]). On every graph $G$, a formula $\delta(\bar{x},\bar{y})$ defines a bipartite graph $G_{\delta}$ on the vertex set $V(G)^{|\bar{x}|}\times V(G)^{|\bar{y}|}$, where two elements $\bar{u}$ and $\bar{v}$ are connected by an edge if the formula $\delta(\bar{u},\bar{v})$ is true in $G$. The semi-ladder index of $\delta$ on $G$ is defined as the semi-ladder index of $G_{\delta}$.
@@
Let $\mathscr{C}$ be a class of graphs and let $\delta$ be a distance formula. Assume that $G_{\delta}$ has bounded semi-ladder index $\ell$ for some fixed $\ell\in\mathbb{N}$ on all graphs $G\in\mathscr{C}$. Then there is an algorithm that solves the domination-type problem $\delta$ on graphs $G$ from $\mathscr{C}$ in time $f(\ell,|\delta|)\cdot\|G\|$.
Annotated Partial Domination can be written with the following first-order formula:
$\varphi_{k,d}=\exists x_{1}\ldots\exists x_{k}$ $\exists y_{1}\ldots\exists y_{d}\forall z\leavevmode\nobreak\ \leavevmode\nobreak\ \Big{(}\bigwedge_{1\leq i\leq k}\neg F(x_{i})\Big{)}\wedge\Big{(}\bigwedge_{1\leq i\leq d}B(y_{i})\Big{)}\wedge$
@@
However, $\varphi$ is not a distance formula as it uses negations. This can be circumvented by inverting the predicates, i.e. $F^{\prime}=V(G)-F$ and $R^{\prime}=V(G)-R$. We then obtain the following formula
@@
While the semi-ladder index is currently one of the most general graph parameters enabling FPT evaluation of $\delta_{k,d}$, some other parameter could have this property.
@@
A graph parameter $t(\cdot)$ has ˜1 if for every graph $G$ Annotated Partial Domination can be solved in time $f(d,k,t(G))\cdot|G|^{\mathcal{O}(1)}$ for a computable function $f$.
As previously discussed, examples of parameters with ˜1 include treewidth, maximum degree, Hadwiger number (i.e. largest clique minor), degeneracy, largest excluded bi-clique, as they all imply bounded semi-ladder index. Other parameters, which are however not as interesting, as the problem can be solved by efficient FO model checking, are cliquewidth and twinwidth (when contraction sequences are given with the input).
As mentioned, we want to solve Annotated Partial Domination on slightly modified graphs. See Definition˜B.6 for the precise definition of bag graphs. We therefore want a parameter that, if bounded on the graph, bounds a parameter with ˜1 on bag graphs.
@@
A graph parameter $t_{2}(\cdot)$ has ˜2 if there is a parameter $t_{1}(\cdot)$ satisfying ˜1 such that for every graph $G$, and integers $q,k,d$, every $(q,k)$-unbreakable tree decomposition $(T,\mathsf{bag})$, vertex $x\in T$, and set $S\subseteq\mathsf{cone}(x)$ we have that $t_{1}(\mathsf{bgraph}_{S}(x))\leq f(q,k,d,t_{2}(G))$.
In the rest of the paper, we precisely reference parameters having ˜1 or ˜2. In particular, the semi-ladder index also satisfies ˜2, see Lemma˜B.11.